OpenStudy (anonymous):

how to change 2x + 3y = 1,200 to slope intercept form?

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

slope-intercept form is y = mx + b where m is the slope and b is the y-intercept

OpenStudy (anonymous):

slope intercept form of line is given as y=mx+c Therefore 2x + 3y = 1,200 can be written as 3y = -2x+1,200 \[y= \frac{-2}{3}x+\frac{1200}{3}\] \[y= \frac{-2}{3}x+400\] which is in the form of y=mx+c where m=slope =-2/3 and c= y-intercept = 400